# OccNeRF: Advancing 3D Occupancy Prediction in LiDAR-Free Environments

### v2

@@ Abstract
As a fundamental task of vision-based perception, 3D occupancy prediction reconstructs 3D structures of surrounding environments. It provides detailed information for autonomous driving planning and navigation. However, most existing methods heavily rely on the LiDAR point clouds to generate occupancy ground truth, which is not available in the vision-based system. In this paper, we propose an OccNeRF method for training occupancy networks without 3D supervision. Different from previous works which consider a bounded scene, we parameterize the reconstructed occupancy fields and reorganize the sampling strategy to align with the cameras’ infinite perceptive range. The neural rendering is adopted to convert occupancy fields to multi-camera depth maps, supervised by multi-frame photometric consistency. Moreover, for semantic occupancy prediction, we design several strategies to polish the prompts and filter the outputs of a pretrained open-vocabulary 2D segmentation model. Extensive experiments for both self-supervised depth estimation and 3D occupancy prediction tasks on nuScenes and SemanticKITTI datasets demonstrate the effectiveness of our method. Code is available at https://github.com/LinShan-Bin/OccNeRF.
Keywords:
3D occupancy prediction LiDAR-free depth estimation1 Introduction
@@
To address this, we propose an OccNeRF method, which targets at training multi-camera occupancy networks without 3D supervision. The overview of our proposed method is shown in Fig. 1. We first utilize a 2D backbone to extract multi-camera 2D features. To save memory, we directly interpolate 2D features to obtain 3D volume features instead of using heavy cross-view attention. In previous works, the volume features are supervised by the bounded occupancy labels (e.g. 50m range) and they only need to predict the occupancy with finite resolution (e.g. 200 200 16). Differently, for LiDAR-free training, we should consider unbounded scenes since the RGB images perceive an infinite range. To this end, we parameterize the occupancy fields to represent unbounded environments. Specifically, we split the whole 3D space into the inside and outside regions. The inside one maintains the original coordinate while the outside one adopts a contracted coordinate. A specific sampling strategy is designed to transfer parameterized occupancy fields to 2D depth maps with neural rendering.
@@
In summary, our key contributions are:
@@
We design a system that trains an occupancy network without LiDAR data, addressing the challenge of sparse surrounding views through the incorporation of temporal information.
@@
We introduce the parameterized occupancy field that allows vision-centric systems to efficiently represent unbounded scenes, matching the extensive perceptual range of cameras.
@@
We devise three strategies to improve the quality of pseudo labels generated by pretrained open-vocabulary segmentation models.
2 Related Work
2.1 3D Occupancy Prediction
@@ 2 Related Work
Due to the significance of the vision-centric autonomous driving systems, more and more researchers begin to focus on 3D occupancy prediction tasks [9, 32, 88, 77, 102, 96, 83, 78, 10, 28, 42, 47, 46]. In the industry community, 3D occupancy is treated as an alternative to LiDAR perception. As one of the pioneering works, MonoScene [9] extracts the voxel features generated by sight projection to reconstruct scenes from a single image. TPVFormer [32] further extends it to multi-camera fashion with tri-perspective view representation. Beyond TPVFormer, SurroundOcc [88] designs a pipeline to generate dense occupancy labels instead of using sparse LiDAR points as the ground truth. In addition, a 2D-3D UNet with cross-view attention layers is proposed to predict dense occupancy. RenderOcc [63] uses the 2D depth maps and semantic labels to train the model, reducing the dependence on expensive 3D occupancy annotations. Compared with these methods, our method does not need any annotated 3D or 2D labels. Recently, as a preprint work, SimpleOccupancy [19] presents a simple while effective framework for occupancy estimation. Although SimpleOccupancy investigates the vision-centric setting, it does not consider the infinite range and semantic prediction.
2.2 Neural Radiance Fields
@@
As one of the most popular topics in 3D area, neural radiance fields (NeRF) [39, 56, 84, 16, 76, 58, 12, 90, 61, 34, 49, 7, 74] have made great achievement in recent years. NeRF [59] learns the geometry of a scene by optimizing a continuous volumetric scene function with multi-view images. To obtain the novel views, volume rendering is performed to convert the radiance fields to RGB images. As a follow-up, mip-NeRF [2] represents the scene at a continuously valued and replaces rays as anti-aliased conical frustums. Beyond mip-NeRF, Zip-NeRF [4] integrates mip-NeRF with a grid-based model for faster training and better quality. There are several extensions of original NeRF, including dynamic scenes [64, 21, 65, 80, 44], model accelerating [82, 91, 93, 62, 86], 3D reconstruction [11, 22, 67, 17, 60, 75], etc. As one of these extensions, some works aim to describe unbounded scenes [95, 3]. NeRF++ [95] split the 3D space as an inner unit sphere and an outer volume and proposes inverted sphere parameterization to represent outside regions. Further, mip-NeRF 360 [3] embeds this idea into mip-NeRF and applies the smooth parameterization to volumes. Inspired by these methods, we also design a parameterization scheme to model the unbounded scene for occupancy prediction.
2.3 Self-supervised Depth Estimation
@@
While early works [38, 98, 18, 50, 69] require dense depth annotations, recent depth estimation methods [25, 85, 55, 92, 81, 66, 72, 71, 37, 79, 99, 14, 6, 101] are designed in a self-supervised manner. Most of these methods predict depth maps and ego-motions simultaneously, adopting the photometric constraints [100, 23] between successive frames as the supervision signal. As a classical work in this field, Monodepth2 [24] proposes some techniques to improve the quality of depth predictions, including the minimum re-projection loss, full-resolution multi-scale sampling, and auto-masking loss. Since modern self-driving vehicles are usually equipped with multiple cameras to capture the surrounding views, researchers begin to concentrate on the multi-camera self-supervised depth estimation task [26, 87, 70, 35, 89, 73]. FSM [26] is the first work to extend monocular depth estimation to full surrounding views by leveraging spatiotemporal contexts and pose consistency constraints. To predict the real-world scale, SurroundDepth [87] uses structure-from-motion to generate scale-aware pseudo depths to pretrain the models. Further, it proposes the cross-view transformer and joint pose estimation to incorporate the multi-camera information. Recently, R3D3 [70] combines the feature correlation with bundle adjustment operators for robust depth and pose estimation. Different from these methods, our approach directly extracts features in 3D space, achieving multi-camera consistency and better reconstruction quality.
3 Approach
3.1 Overview
@@
Fig. 2 shows the pipeline of our approach. With the multi-camera images as inputs, we first utilize a 2D backbone to extract cameras’ features . Then the 2D features are interpolated to the 3D space to obtain the volume features with known intrinsic and extrinsic . As discussed in Section 3.2, to represent the unbounded scenes, we propose a coordinate parameterization to contract the infinite range to a limited occupancy field. The volume rendering is performed to convert occupancy fields to multi-frame depth maps, which are supervised by photometric loss. Section 3.3 introduces this part in detail. Finally, Section 3.4 shows how we use a pretrained open-vocabulary segmentation model to get 2D semantic labels.
3.2 Parameterized Occupancy Fields
@@
where is the normalized coordinate of the input and indicates the normalized parameterized coordinate. is the bound of the inside region, which is different for x, y, z directions. represents the proportion of the region of interest in the parameterized space. Higher indicates we use more space to describe the inside region. Note that the two functions in Equation 1 have the same value and gradient at . Please refer to the supplementary material for the derivation details.
@@
where proj is the function projecting 3D points to the 2D image plane defined by the camera extrinsic and intrinsic , is the bilinear interpolation operator, is the interpolation result. To simplify the aggregation process and reduce computation costs, we directly average the multi-camera 2D features to get volume features, which is the same as the method used in [19, 27]. Finally, a 3D convolution network [20] is employed to extract features and predict final occupancy outputs.
3.3 Multi-frame Depth Estimation
To project the occupancy fields to multi-camera depth maps, we adopt volume rendering [57], which is widely used in NeRF-based methods [59, 95, 2]. To render the depth value of a given pixel, we cast a ray from the camera center along the direction pointing to the pixel. The ray is represented by . Then, we sample points along the ray in 3D space to get the density . For the selected quadrature points, the depth of the corresponding pixel is computed by:
@@ 3.4 Semantic Supervision
Specifically, when dealing with categories, we employ three strategies to determine the prompts provided to the Grounding DINO. These strategies consist of synonymous substitution, where we replace words with their synonyms (e.g., changing ‘car’ to ‘sedan’ to enable the model to distinguish it from ‘truck’ and ‘bus’); splitting single words into multiple entities (e.g., ‘manmade’ is divided into ‘building’, ‘billboard’, and ‘bridge’ etc. to enhance differentiation); and incorporating additional information (e.g., introducing ‘bicyclist’ to facilitate the detection of a person on a bike). Subsequently, we obtain detection bounding boxes along with their corresponding logits and phrases, which are fed to SAM [36] to generate precise segmentation binary masks. After multiplying the Grounding DINO logits with binary masks, every pixel has logits. We get the per-pixel label using:
@@
To leverage the 2D semantic supervision, we initially utilize a semantic head with output channels to map volume features extracted to semantic outputs, denoted as . Similar to the method outlined in Section 3.3, we engage in volume rendering once more using the subsequent equation:
@@
4 Experiments
4.1 Experimental Setup
4.2 Self-supervised Depth Estimation
@@
| SimpleOcc [19] | 7.99 | - | 0.67 | 1.18 | 3.21 | 7.63 | 1.02 | 0.26 | 1.80 | 0.26 | 1.07 | 2.81 | 40.44 | 18.30 | 17.01 | 13.42 | 10.84 | |
| OccNeRF | 10.81 | - | 0.83 | 0.82 | 5.13 | 12.49 | 3.50 | 0.23 | 3.10 | 1.84 | 0.52 | 3.90 | 52.62 | 20.81 | 24.75 | 18.45 | 13.19 |
Evaluation Metric: For depth estimation, we use the commonly used depth evaluation metrics [24, 100, 87]: Abs Rel, Sq Rel, RMSE, RMSE log and . The Abs Rel is the main metric and see supplementary for the details of these metrics. During evaluation, we do not perform median scaling since our method can predict real-world scale. For 3D occupancy prediction, we use the mean intersection over union (mIoU) of all classes as the semantic-aware metric and intersection over union (IoU), precision and recall as semantic-agnostic metrics.
Table 1 shows the self-supervised multi-camera depth estimation results on nuScenes dataset. We do not use pretrained segmentation models in this experiment. The results are averaged over 6 cameras and ‘FSM*’ is the reproduced FSM [26] result reported in [35]. We can see that our method outperforms other SOTA methods by a large margin, demonstrating the effectiveness of OccNeRF. Compared with depth estimation methods, our method directly predicts occupancy in 3D space, naturally guaranteeing multi-camera consistency. Moreover, we do not need to lift 2D depth maps to 3D point clouds with post-processing.
@@
4.3 Semantic Occupancy Prediction
We conduct experiments on semantic occupancy prediction using the Occ3D-nuScenes dataset. The pretrained open-vocabulary model [36, 51] struggles with ambiguous prompts like ’other’ and ’other flat’, so we exclude them during the evaluation. To compare with the SimpleOcc method [19], we add a semantic head to the original model and leverage the generated 2D semantic labels to train it. Our approach significantly outperforms SimpleOcc, as detailed in Table 2, and achieves competitive results against some fully-supervised methods. Notably, it excels in predicting ’drivable space’ and ’manmade’ classes, outdoing all supervised methods. However, it falls short in detecting small objects, such as bicycles and pedestrians, where it lags behind state-of-the-art supervised methods, likely due to the open-vocabulary model’s limitations in capturing small details.
We further explored geometry occupancy prediction with the nuScenes [8] and SemanticKITTI [5]. Since most works reported in occ3D-nuScenes [77] do not provide codes, we can only evaluate RenderOcc [63] and SeimpleOcc [19]. As detailed in Table 3 and Table 7, our approach significantly outperforms other methods supervised by images and achieves competitive results against methods with even stronger supervision.

### v3

@@ Abstract
Occupancy prediction reconstructs 3D structures of surrounding environments. It provides detailed information for autonomous driving planning and navigation. However, most existing methods heavily rely on the LiDAR point clouds to generate occupancy ground truth, which is not available in the vision-based system. In this paper, we propose an OccNeRF method for training occupancy networks without 3D supervision. Different from previous works which consider a bounded scene, we parameterize the reconstructed occupancy fields and reorganize the sampling strategy to align with the cameras’ infinite perceptive range. The neural rendering is adopted to convert occupancy fields to multi-camera depth maps, supervised by multi-frame photometric consistency. Moreover, for semantic occupancy prediction, we design several strategies to polish the prompts and filter the outputs of a pretrained open-vocabulary 2D segmentation model. Extensive experiments for both self-supervised depth estimation and 3D occupancy prediction tasks on nuScenes and SemanticKITTI datasets demonstrate the effectiveness of our method. The code is available at https://github.com/LinShan-Bin/OccNeRF.
Index Terms:
3D occupancy prediction, LiDAR-free, self-supervised depth estimationI Introduction
@@
To address this, we propose an OccNeRF method, which targets at training multi-camera occupancy networks without 3D supervision. The overview of our proposed method is shown in Fig. 1. We first utilize a 2D backbone to extract multi-camera 2D features. To save memory, we directly interpolate 2D features to obtain 3D volume features instead of using heavy cross-view attention. In previous works, the volume features are supervised by the bounded occupancy labels (e.g. 50m range) and they only need to predict the occupancy with finite resolution (e.g. 200 200 16). Differently, for LiDAR-free training, we should consider unbounded scenes since the RGB images perceive an infinite range. To this end, we parameterize the occupancy fields to represent unbounded environments. Specifically, we split the whole 3D space into the inside and outside regions. The inside one maintains the original coordinate while the outside one adopts a contracted coordinate. A specific sampling strategy is designed to transfer parameterized occupancy fields to 2D depth maps with the neural rendering algorithm.
@@
In summary, our principal contributions include:
@@
We develop a system that trains an occupancy network without the need for LiDAR data, addressing the challenge of sparse surrounding views by integrating temporal information for more geometry information.
@@
We introduce a parameterized occupancy field that enables vision-centric systems to efficiently represent unbounded scenes, aligning with the extensive perceptual capabilities of the cameras.
@@
We devise a pipeline to generate high-quality pseudo labels with pretrained open-vocabulary segmentation models, with three prompt strategies to improve the accuracy.
II Related Work
This section examines three interrelated areas in computer vision: 3D occupancy prediction, neural radiance fields, and self-supervised depth estimation. We highlight key advancements and ongoing challenges, providing a critical overview that identifies gaps in current research and suggests avenues for further investigation.
II-A 3D Occupancy Prediction
@@ II Related Work
Due to the significance of the vision-centric autonomous driving systems, more and more researchers begin to focus on 3D occupancy prediction tasks [16, 22, 3, 13, 23, 14, 15, 24, 25, 26, 27, 28, 29]. In the industry community, 3D occupancy is treated as an alternative to LiDAR perception. As one of the pioneering works, MonoScene [16] extracts the voxel features generated by sight projection to reconstruct scenes from a single image. TPVFormer [22] further extends it to multi-camera fashion with tri-perspective view representation. Beyond TPVFormer, SurroundOcc [3] designs a pipeline to generate dense occupancy labels instead of using sparse LiDAR points as the ground truth. In addition, a 2D-3D UNet with cross-view attention layers is proposed to predict dense occupancy. RenderOcc [30] uses the 2D depth maps and semantic labels to train the model, reducing the dependence on expensive 3D occupancy annotations. Compared with these methods, our method does not need any annotated 3D or 2D labels. Occ3D [13] establishes the occupancy benchmarks used in CVPR 2023 occupancy prediction challenge and proposes a coarse-to-fine occupancy network. SimpleOccupancy [31] presents a simple while effective framework for occupancy estimation. Although SimpleOccupancy [31] and SelfOcc [selfocc] investigate the vision-centric setting, they do not consider the infinite perception range of cameras.
II-B Neural Radiance Fields
@@
As one of the most popular topics in 3D area, neural radiance fields (NeRF) [17] have made great achievement in recent years. NeRF [17] learns the geometry of a scene by optimizing a continuous volumetric scene function with multi-view images. To obtain the novel views, volume rendering is performed to convert the radiance fields to RGB images. As a follow-up, mip-NeRF [32] represents the scene at a continuously valued and replaces rays as anti-aliased conical frustums. Beyond mip-NeRF, Zip-NeRF [33] integrates mip-NeRF with a grid-based model for faster training and better quality. There are several extensions of original NeRF, including dynamic scenes [34, 35, 36, 37, 38, 39, 40], 3D reconstruction [41, 42, 43, 44, 45], model accelerating [46, 47, 48, 49, 50, 51, 52, 53], etc. As one of these extensions, some works aim to describe unbounded scenes [54, 55]. NeRF++ [54] split the 3D space as an inner unit sphere and an outer volume and proposes inverted sphere parameterization to represent outside regions. Further, mip-NeRF 360 [55] embeds this idea into mip-NeRF and applies the smooth parameterization to volumes. Inspired by these methods, we also design a parameterization scheme to model unbounded scenes for the occupancy prediction task.
II-C Self-supervised Depth Estimation
@@
While early works [56, 57, 58, 59, 60] require dense depth annotations, recent depth estimation methods [61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73] are designed in a self-supervised manner. Most of these methods predict depth maps and ego-motions simultaneously, adopting the photometric constraints [74, 75] between successive frames as the supervision signal. As a classical work in this field, Monodepth2 [76] proposes some techniques to improve the quality of depth predictions, including the minimum re-projection loss, full-resolution multi-scale sampling, and auto-masking loss. Since modern self-driving vehicles are usually equipped with multiple cameras to capture the surrounding views, researchers begin to concentrate on the multi-camera self-supervised depth estimation task [77, 78, 79, 80, 81, 82]. FSM [77] is the first work to extend monocular depth estimation to full surrounding views by leveraging spatiotemporal contexts and pose consistency constraints. To predict the real-world scale, SurroundDepth [78] uses structure-from-motion to generate scale-aware pseudo depths to pretrain the models. Further, it proposes the cross-view transformer and joint pose estimation to incorporate the multi-camera information. Recently, R3D3 [79] combines the feature correlation with bundle adjustment operators for robust depth and pose estimation. Different from these methods, our approach directly extracts features in 3D space, achieving multi-camera consistency and better reconstruction quality.
III Method
@@
Fig. 2 shows the pipeline of our approach. With the multi-camera images as inputs, we first utilize a 2D backbone to extract cameras’ features . Then the 2D features are interpolated to the 3D space to obtain the volume features with known intrinsic and extrinsic . As discussed in Section III-A, to represent the unbounded scenes, we propose a coordinate parameterization to contract the infinite range to a limited occupancy field. The volume rendering is performed to convert occupancy fields to multi-frame depth maps, which are supervised by photometric loss. Section III-B introduces this part in detail. Finally, Section III-C shows how we use a pretrained open-vocabulary segmentation model to get 2D semantic labels.
III-A Parameterized Occupancy Fields
@@
where represents the proportion of the region of interest in the parameterized space. Higher indicates we use more space to describe the inside region. denotes the normalized coordinate based on the input and pre-defined inside region bound . The parameters and are introduced to maintain the continuity of the first derivative. The determination of these parameters is achieved through the resolution of the ensuing equations:
The derived solutions are presented as:
@@
where proj is the function projecting 3D points to the 2D image plane defined by the camera extrinsic and intrinsic , is the bilinear interpolation operator, is the interpolation result. To simplify the aggregation process and reduce computation costs, we directly average the multi-camera 2D features to get volume features, which is the same as the method used in [31, 83]. Finally, a 3D convolution network [84] is employed to extract features and predict the final occupancy outputs.
III-B Multi-frame Depth Estimation
To project the occupancy fields to multi-camera depth maps, we adopt volume rendering [85], which is widely used in NeRF-based methods [17, 54, 32]. To render the depth value of a given pixel, we cast a ray from the camera center o along the direction d pointing to the pixel. The ray is represented by . Then, we sample points along the ray in 3D space to get the density . For the selected quadrature points, the depth of the corresponding pixel is computed by:
@@ III-C Semantic Supervision
Specifically, when dealing with categories, we employ three strategies to determine the prompts provided to the Grounding DINO. These strategies consist of synonymous substitution, where we replace words with their synonyms (e.g., changing ‘car’ to ‘sedan’ to enable the model to distinguish it from ‘truck’ and ‘bus’); splitting single words into multiple entities (e.g., ‘manmade’ is divided into ‘building’, ‘billboard’, and ‘bridge’ etc. to enhance differentiation); and incorporating additional information (e.g., introducing ‘bicyclist’ to facilitate the detection of a person on a bike). See TABLE I for more details. Subsequently, we obtain detection bounding boxes along with their corresponding logits and phrases, which are fed to SAM [18] to generate precise segmentation binary masks. After multiplying the Grounding DINO logits with binary masks, every pixel has logits. We get the per-pixel label using:
@@
To leverage the 2D semantic supervision, we initially utilize a semantic head with output channels to map volume features extracted to semantic outputs, denoted as . Similar to the method outlined in Section III-B, we engage in volume rendering once more using the subsequent equation:
@@
| Original labels | Ours |
| car | sedan |
| bicycle | bicycle bicyclist |
| vegetation | tree |
| motorcycle | motorcycle motorcyclist |
| drivable surface | highway |
| traffic cone | cone |
| construction vehicle | crane |
| manmade | building compound bridge |
| pole billboard light ashbin |
@@
| front left | front | front right | back right | back | back left |
@@
| SimpleOcc [31] | 7.99 | 7.05 | 0.67 | 1.18 | 3.21 | 7.63 | 1.02 | 0.26 | 1.80 | 0.26 | 1.07 | 2.81 | 40.44 | 18.30 | 17.01 | 13.42 | 10.84 | |
| OccNeRF | 10.81 | 9.53 | 0.83 | 0.82 | 5.13 | 12.49 | 3.50 | 0.23 | 3.10 | 1.84 | 0.52 | 3.90 | 52.62 | 20.81 | 24.75 | 18.45 | 13.19 |
| front left | front | front right | back right | back | back left |
IV Experiment
IV-A Experimental Setup
IV-B Self-supervised Depth Estimation
Evaluation Metric: For depth estimation, we use the commonly used depth evaluation metrics [76, 74, 78] as outlined in the following:
Abs Rel: ,
Sq Rel: ,
RMSE: ,
RMSE log: ,
: % of s.t. ,
where and indicate predicted and ground truth depths respectively, and T indicates all pixels on the depth image . In our experiments, all the predicted depth maps are scale-aware and we do not perform any scale alignment. The Abs Rel is the main metric for depth estimation tasks and it reveals the relative errors of estimated depths. Note that during evaluation we do not perform median scaling since our method can predict real-world scale given ground truth poses. For the 3D occupancy predictiontask, we use the mean intersection over union (mIoU) of all classes as the semantic-aware metric and the intersection over union (IoU), precision and recall as the semantic-agnostic metrics.
TABLE II shows the self-supervised multi-camera depth estimation results on nuScenes dataset. We do not use pretrained segmentation models in this experiment. The results are averaged over 6 cameras and ‘FSM*’ is the reproduced FSM [77] result reported in [80]. We can see that our method outperforms other SOTA methods by a large margin, demonstrating the effectiveness of OccNeRF. Compared to previous depth estimation methods, our approach directly predicts an occupancy field in 3D space instead of predicting per-pixel depths. This naturally guarantees multi-camera consistency. Furthermore, our method eliminates the need for post-processing steps that lift 2D depths to 3D point clouds.
@@ IV-B Self-supervised Depth Estimation
IV-C Occupancy Prediction
We conduct experiments on semantic occupancy prediction using the Occ3D-nuScenes dataset. The pretrained open-vocabulary model [18, 19] struggles with ambiguous prompts like ‘other’ and ‘other flat’, so we exclude them during the evaluation. To compare with the SimpleOcc method [31], we add a semantic head to the original model and leverage the generated 2D semantic labels to train it. Our approach significantly outperforms SimpleOcc, as detailed in TABLE III, and achieves competitive results against some fully-supervised methods. Notably, it excels in predicting ‘drivable space’ and ‘manmade’ classes, outdoing all supervised methods. However, it falls short in detecting small objects, such as bicycles and pedestrians, where it lags behind state-of-the-art supervised methods, likely due to the open-vocabulary model’s limitations in capturing small details.
We further explored the geometry occupancy prediction task with the nuScenes [20] and SemanticKITTI [21]. Since most works reported in occ3D-nuScenes [13] do not provide codes, we can only evaluate RenderOcc [30] and SeimpleOcc [31]. As detailed in TABLE IV and TABLE V, our approach outperforms other methods supervised by images and achieves competitive results against methods with stronger supervision.